• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Xinhu Yan

    Log entry time 16:05:51 on December 7,2007

    Entry number 218863

    keyword=Day shift Summary

    Day Shift Summary
    SL:Xinhu Yan
    TO:Joaqin Lopez Herraiz
    3rd:Jose Udias

    8:15: Both of spectrometer have been moved to 15deg
    8:23 Call Mcc taking beam back but have some problem about water flow diffuser(?) in the Hall
    8:45 Call MCC for control access for checking the water diffuser problem in the Hall
    10:41 the diffuser Water flow problem fixed .we ask MCC for beam back
    10:59 beam come back ,we start to take production data
    11:51 the rate for Pb target is so high about 700KHz even at 5uA, JP suggest to adjust prescale to reduce the rate to about 1KHz
    12:13 Alexandre was changing the High Voltage for the NaI detector
    12:43 MCC call that beam would be away for about 15 minutes.
    12:59 changed the Momentum of left arm to 983.473Mev/C for checking High trigger rate and Momentum of right arm to the next setting 1131.327Mev/C while waiting for the beam
    14:02 beam come back and alexandre still working on the NaI HV
    15:21 recycle the left arm Q2/Q3 with 1600A and change the momentum of right arm to 1054.81Mev/c
    15:25 the recycle of left arm was done and change the momentum of left arm to 1131.32Mev/c

    --------------------------------------------------------------
    Left Spectrometer
    E=1256Mev 15deg

    P0=1213.39Mev/C
    Run 3450 target Fe left 10:58-11:03 evt.1.26M
    Run 3451 target C12 left 11:06-11:12 evt.1.29M
    Run 3452 target dummy 11:18-11:23 evt.1.29M
    Run 3453 target Pb 11:39-11:45 evt.302K
    Run 3454 target H2 11:53-12:00 evt.1.6M
    Run 3455 target He 12:08-12:13 evt.1.5M

    P0=983.473Mev/c
    Run 3456 Cosmic
    Run 3457 target He 14:02-14:07 evt.419K
    Run 3458 target Pb 14:19-14:23 evt.820K
    Run 3459 target H2 14:30-14:38 evt.1.1M
    Run 3460 target Dummy 14:45-14:53 evt.516K
    Run 3461 target C12 15:03-15:08 evt.207K
    Run 3462 target Fe 15:12-15:19 evt.948K

    P0=1131.327Mev/C
    Run 3463 target Fe 15:34-15:39 evt.1.2M
    Run 3464 target C12 15:43-15:49 evt.1.2M
    Run 3465 3466 target Dummy Junk runs
    Run 3467 target Dummy 15:59-16:05 evt.925K

    ---------------------------------------------------------------
    Right Spectrometer
    E=1256Mev 15deg

    P0=1213.39Mev/C
    Run 21563 target Fe left 10:58-11:03 evt.1.29M
    Run 21564 target C12 left 11:06-11:12 evt.1.28M
    Run 21565 target dummy 11:18-11:33 evt.1.29M
    Run 21566 target Pb 11:40-11:45 evt 291K
    Run 21567 target H2 11:53-12:00 evt.1.6M
    Run 21568 target He 12:07-12:13 evt.1.5M

    P0=1131.327Mev/c
    Run 21568 Cosmic
    Run 21569 target He 14:04-14:09 evt.830K
    Run 21570 target Pb 14:15-14:23 evt.1.3M
    Run 21571 target H2 14:29-14:38 evt.1.3M
    Run 21572 target Dummy 14:44-14:53 evt.557K
    Run 21573 target C12 15:00-15:08 evt.284K
    Run 21574 target Fe 15:12-15:19 evt.925K